How do I find the vaule of x to this proplem
JKLM,MK =8x-9, LM=10x+2,and JL=7x+10

Answer by Timnewman(323)

Hi dear,
If you try to draw the shape formed by JKLM,you will notice that JL is parallel and equal to MK.
SINCE JL=MK
therefore,7x+10=8x-9
collect like-terms
8x-7x=10+9
x=19
The value of x is 19
HOPE THIS HELPS,..